# CHASE: 3D-Consistent Human Avatars with Sparse Inputs via Gaussian Splatting and Contrastive Learning

### v2

@@ Abstract
Recent advancements in human avatar synthesis have utilized radiance fields to reconstruct photo-realistic animatable human avatars. However, both NeRFs-based and 3DGS-based methods struggle with maintaining 3D consistency and exhibit suboptimal detail reconstruction, especially with sparse inputs. To address this challenge, we propose CHASE, which introduces supervision from intrinsic 3D consistency across poses and 3D geometry contrastive learning, achieving performance comparable with sparse inputs to that with full inputs. Following previous work, we first integrate a skeleton-driven rigid deformation and a non-rigid cloth dynamics deformation to coordinate the movements of individual Gaussians during animation, reconstructing basic avatar with coarse 3D consistency. To improve 3D consistency under sparse inputs, we design Dynamic Avatar Adjustment(DAA) to adjust deformed Gaussians based on a selected similar pose/image from the dataset. Minimizing the difference between the image rendered by adjusted Gaussians and the image with the similar pose serves as an additional form of supervision for avatar. Furthermore, we propose a 3D geometry contrastive learning strategy to maintain the 3D global consistency of generated avatars. Though CHASE is designed for sparse inputs, it surprisingly outperforms current SOTA methods in both full and sparse settings on the ZJU-MoCap and H36M datasets, demonstrating that our CHASE successfully maintains avatar’s 3D consistency, hence improving rendering quality. The code will be made available.
@@
footnotetext: Haoyu Zhao completed this work during an internship at Shanghai Jiao Tong University.
Introduction
@@
Photorealistic rendering and animation of human bodies is a crucial research area with wide-ranging applications in AR/VR, visual effects, visual try-on, and movie production (Zackariasson and Wilson 2012; Healey, Wang, and et al 2021). Early works (Mildenhall et al. 2021; Oechsle, Peng, and Geiger 2021; Niemeyer et al. 2020; Peng et al. 2021b; Wang et al. 2022a; Sitzmann et al. 2021) for creating human avatars relied on capturing high-quality data through multi-camera setups, which is extensive computing and needs lots of manual effort. While these methods excel at creating a single scene/object from sufficient input views, it is very challenging for them to generalize to new scenes/objects with few samples (Kwon et al. 2024).
@@
Recent advancements have explored using neural radiance fields (NeRF) for modeling 3D human avatars (Mildenhall et al. 2021; Wang et al. 2022a), typically employing parametric body models to model deformations. Some methods (Chen et al. 2022; Zhao et al. 2022) use human template models to facilitate generalizable and robust synthesis. Though NeRF-based methods have made significant progress in generalizable human rendering, they are less efficient to train and render due to their computationally intensive per-pixel volume rendering process.
@@
Point-based rendering (Zheng et al. 2023) has emerged as an efficient alternative to NeRFs, offering significantly faster rendering. The recently proposed 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) gains popularity for its efficient rasterization-based rendering speed. Numerous works have further explored the 3D Gaussian representation for dynamic 3D human avatars (Lei et al. 2024; Moreau et al. 2024; Shao et al. 2024; Hu et al. 2024b; Kocabas et al. 2024; Wang et al. 2024a; Qian et al. 2024; Hu et al. 2024a). However, these methods often face challenges in maintaining 3D consistency and producing high-quality reconstructions, particularly with sparse inputs.
@@
To address the aforementioned issues, we propose CHASE, which is capable of reconstructing 3D Consistent Human Avatars with Sparse inputs via Gaussian Splatting and contrastivE learning. We first integrates a skeleton-driven rigid deformation, and a non-rigid cloth dynamics deformation to create human avatar with coarse 3D consistency. To enhance 3D consistency under sparse inputs, we utilize the intrinsic 3D consistency of images across different poses within the same person. Specifically, for each training pose/image, we select a similar pose/image from the dataset and then adjust the deformed Gaussians using the proposed Dynamic Avatar Adjustment (DAA), an explicit point-based control graph adjustment strategy, to the selected similar pose. Then we minimize differences between the rendered image of the adjusted Gaussians and the image corresponding to the selected similar pose, which serves as an additional form of supervision for human avatar. Additionally, we employ 3D geometry contrastive learning, utilizing features from a 3D feature extractor, to further enhance the global 3D consistency of generated human avatars. We conduct extensive experiments on the ZJU-MoCap data (Peng et al. 2021c), H36M (Ionescu et al. 2013) and find that CHASE not only achieves better performance in sparse inputs setting but also outperforms other SOTAs in full setting.
In summary, our work makes the following contributions:
@@
  • We propose 3D geometry contrastive learning to enforce consistency across different representations of the same pose and facilitate effective 3D global understanding.
@@
Related Work
Contrastive Representation Learning
@@
Contrastive Representation Learning is one of the mainstream self-supervised learning paradigms (Hadsell, Chopra, and LeCun 2006), which learns potential semantics from constructed invariance or equivariance (Rumen et al. 2022). In 3D, PointContrast (Xie et al. 2020) proposes geometric augmentation to generate positive and negative pairs. CrossPoint (Afham et al. 2022) uses both inter- and intra-modal contrastive learning. PointCLIP (Zhang et al. 2022) achieves image-point alignment by projecting point clouds onto 2D depth images. RECON (Qi et al. 2023) focuses on single- and cross-modal contrastive learning through discriminative contrast (Khosla et al. 2020) or global feature alignment (Radford et al. 2021). Our CHASE introduces a 3D geometry contrastive learning method across different poses to enforce consistency across different representations of the same pose.
3D Editing and Deformation
@@
Traditional deformation methods in computer graphics are typically based on Laplacian coordinates (Gao et al. 2019; Sorkine and Alexa 2007; Sorkine 2005), Poisson equations (Yu et al. 2004), and cage-based methods (Yifan et al. 2020). However, these methods often rely on implicit and computationally expensive NeRF-based approaches.
@@
Numerous works (Chen et al. 2024; Wang et al. 2024b) have proposed techniques for editing 3D Gaussian Splatting (3DGS)(Kerbl et al. 2023). SuGaR(Guédon and Lepetit 2024) introduces a mesh extraction method that produces meshes from 3DGS, which can then be edited. SC-GS (Huang et al. 2024) proposes deforming Gaussians by transferring the movement of control points. Our CHASE employs a novel explicit point-based control graph deformation strategy, which is more intuitive and efficient.
3D Human Modeling
@@
Since the high-quality rendering achieved by the seminal work Neural Radiance Fields (NeRF)(Mildenhall et al. 2021), there has been a surge of research on neural rendering for human avatars(Liu et al. 2021; Li et al. 2023; Peng et al. 2021c). Although NeRF is designed for static objects, HumanNeRF (Weng et al. 2022) extends NeRF to enable capturing dynamic human motion using just a single monocular video. Neural Body (Peng et al. 2021c) associates a latent code with each SMPL (Loper et al. 2015) vertex to encode appearance, which is then transformed into observation space based on the human pose. Furthermore, Neural Actor (Liu et al. 2021) learns a deformable radiance field with SMPL (Loper et al. 2015) as guidance and utilizes a texture map to improve the final rendering quality. Posevocab (Li et al. 2023) designs joint-structured pose embeddings to encode dynamic appearances under different key poses, allowing for more effective learning of joint-related appearances. However, a major limitation of NeRF-based methods is that NeRFs are slow to train and render.
@@
Point-based rendering (Zheng et al. 2023) has proven to be an efficient alternative to NeRFs for fast inference and training. Extending point clouds to 3D Gaussians, 3D Gaussian Splatting (3DGS)(Kerbl et al. 2023) models the rendering process by splatting a set of 3D Gaussians onto the image plane via alpha blending. Given the impressive performance of 3DGS in both quality and speed, numerous works have further explored the 3D Gaussian representation for dynamic 3D human avatar reconstruction(Lei et al. 2024; Qian et al. 2024; Hu et al. 2024b; Kocabas et al. 2024). Human Gaussian Splatting (Moreau et al. 2024) showcases 3DGS as an efficient alternative to NeRF. SplattingAvatar (Shao et al. 2024) and GomAvatar (Wen et al. 2024) extend lifted optimization to simultaneously optimize the parameters of the Gaussians while walking on the triangle mesh. However, these methods struggle to maintain 3D consistency and produce low-quality reconstructions when applied to human avatar creation with only sparse inputs. Our CHASE introduces a novel 2D image supervision to 3D human body modeling and 3D geometry contrastive learning, enhancing the 3D consistency of human avatars.
@@
Preliminaries
SMPL (Loper et al. 2015). The SMPL model is a pre-trained parametric human model representing body shape and pose. In SMPL, body shape and pose are controlled by pose and shape. In this work, we apply the Linear Blend Skinning (LBS) algorithm used in SMPL to transform points from a canonical space to a posed space.
LBS (Sumner, Schmid, and Pauly 2007). Linear Blend Skinning (LBS) is a weight-based technique that associates each vertex with one or more joints and uses weight values to describe the influence of each joint on the vertex. Vertex deformation is calculated by linearly interpolating transformations on the associated joints: $\mathcal{X}_{v}^{{}^{\prime}}=\sum_{j=1}^{J}w_{j}(\mathcal{X}_{v})B_{j}\mathcal{X}_{v}$, where $J$ represents the number of joints, $N$ represents the number of vertices, $\mathcal{X}_{v}^{{}^{\prime}}\in\mathbb{R}^{N\times 3}$ is the new position of the skinned vertex, $w\in\mathbb{R}^{N\times J}$ is the skinning weight matrix, $B\in\mathbb{R}^{J\times 4\times 4}$ is the affine transformation matrix of each joint representing rotation and translation, (i.e. bone transforms) and $\mathcal{X}_{v}\in\mathbb{R}^{N\times 3}$ is the original mesh vertex position.
3D Gaussian Splatting (3DGS) (Kerbl et al. 2023). 3DGS explicitly represents scenes using point clouds, where each point is modeled as a 3D Gaussian defined by a covariance matrix $\Sigma$ and a center point $\mathcal{X}$, the latter referred to as the mean. The value at point $\mathcal{X}$ is:
@@
In rendering novel views, differential splatting as introduced by (Yifan et al. 2019), involves using a viewing transform $W$ and the Jacobian matrix $J$ of the affine approximation of the projective transformation to compute the transformed covariance matrix: $\Sigma^{\prime}=JW\Sigma W^{T}J^{T}$. The color and opacity at each pixel are computed from the Gaussian’s representations: $G(\mathcal{X})=e^{-\frac{1}{2}\mathcal{X}^{T}\Sigma^{-1}\mathcal{X}}.$ The blending of $N$ ordered points overlapping a pixel is given by the formula:
$\mathcal{C}=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{i}),$ (2)
where $c_{i}$, $\alpha_{i}$ represent the density and color of this point computed by a 3D Gaussian $G$ with covariance $\Sigma$ multiplied by an optimizable per-point opacity and SH color coefficients.
Method
@@
Non-rigid and Rigid Deformation
Inspired by (Weng et al. 2022; Qian et al. 2024), we deform 3D Gaussians from canonical space ${\mathcal{G}_{c}}$ to observation space ${\mathcal{G}_{o}}$ by integrating a rigid articulation with a non-rigid transformation. We employ a non-rigid deformation network that takes the canonical positions $\mathcal{X}_{c}$ of the 3D Gaussians ${\mathcal{G}c}$ and a pose latent code which encodes SMPL pose $p_{i}$ using a lightweight hierarchical pose encoder (Mihajlovic et al. 2021) as input. The network then outputs the offsets for various parameters of the 3D Gaussians $\mathcal{G}_{c}$: $\Delta{(\mathcal{X},\mathcal{C},\alpha,s,r)}$. The canonical Gaussians are deformed by:
@@
Figure 2: For each training pose/image, we select similar poses/images from the dataset and then adjust the deformed Gaussians using Dynamic Avatar Adjustment (DAA). By minimizing the discrepancy between the rendered image of the adjusted avatar and the selected similar pose image, we introduce an additional supervision, thereby enhancing the creation of animatable avatars.
Dynamic Avatar Adjustment
To address extremely sparse inputs, we leverage the intrinsic 3D consistency of human avatars across different poses/images, as shown in Fig. 2. Specifically, for each training pose/image, we select a similar pose $p^{a}_{i}$ with its paired image $x_{i}^{a}$ from the dataset and then use a dense motion field $F_{adj}$ as an additional adjustment to transform deformed Gaussians ${\mathcal{G}_{o}}$ into adjusted Gaussians ${\mathcal{G}_{a}}$, aligning them with the selected pose/image ($p_{i}^{a}$/$x_{i}^{a}$). In this way, we successfully introduce an additional 2D image supervision, improving the 3D consistency of human avatars.
To achieve precise control of the 3D Gaussians, we sample 6,890 points from the SMPL model (Loper et al. 2015) as our sparse control points in canonical space. Then, we obtain the dense motion field using LBS by locally inheriting the LBS weights from neighboring control points. Specifically, for each 3D Gaussian, we use the k-nearest neighbor (KNN) search to find its nearest neighboring control points in canonical space. The entire adjustment process is as:

### v3

@@ Abstract
Existing approaches for human avatar generation–both NeRF-based and 3D Gaussian Splatting (3DGS) based–struggle with maintaining 3D consistency and exhibit degraded detail reconstruction, particularly when training with sparse inputs. To address this challenge, we propose CHASE, a novel framework that achieves dense-input-level performance using only sparse inputs through two key innovations: cross-pose intrinsic 3D consistency supervision and 3D geometry contrastive learning. Building upon prior skeleton-driven approaches that combine rigid deformation with non-rigid cloth dynamics, we first establish baseline avatars with fundamental 3D consistency. To enhance 3D consistency under sparse inputs, we introduce a Dynamic Avatar Adjustment (DAA) module, which refines deformed Gaussians by leveraging similar poses from the training set. By minimizing the rendering discrepancy between adjusted Gaussians and reference poses, DAA provides additional supervision for avatar reconstruction. We further maintain global 3D consistency through a novel geometry-aware contrastive learning strategy. While designed for sparse inputs, CHASE surpasses state-of-the-art methods across both full and sparse settings on ZJU-MoCap and H36M datasets, demonstrating that our enhanced 3D consistency leads to superior rendering quality. Project page: https://chaseprojectpage.github.io/.
Figure 1: CHASE. We propose an efficient method for creating 3D-consistent animatable avatars from just videos. Our method achieve better quality to the most recent SOTA methods [37, 8, 29] in both full and sparse inputs.
@@
footnotetext: * Equal contributions.footnotetext: †Corresponding Author.footnotetext: Haoyu Zhao completed this work during an internship at Shanghai Jiao Tong University.
1 Introduction
@@
Photo-realistic rendering and animation of human bodies is a critical research area with wide-ranging applications in AR/VR, visual effects, virtual try-on, and film production [6]. Early approaches [23] relied on multi-camera setups to capture high-quality data, requiring extensive computational resources and manual effort. While these methods perform well for reconstructing a single scene or object with sufficient input views, they struggle to generalize to new scenes or objects from limited samples [15].
@@
Recent advancements have explored using neural radiance fields (NeRF) for modeling 3D human avatars [21], typically employing parametric body models to model deformations. Some methods [2, 46, 11] use human template models to facilitate generalizable and robust synthesis. However, NeRF-based methods are less efficient to train and render due to their computationally intensive per-pixel volume rendering process.
@@
Point-based rendering [48] has emerged as an efficient alternative to NeRFs, offering significantly faster rendering. The recently proposed 3D Gaussian Splatting (3DGS) [12] gains popularity for its fast rendering speed. Numerous works have further explored the 3D Gaussian representation for dynamic 3D human avatars [16, 22, 32, 8, 14, 34, 29, 7]. However, these methods often face challenges in maintaining 3D consistency and producing high-quality reconstructions, particularly with sparse inputs.
@@
To address the aforementioned issues, we propose CHASE, which is capable of reconstructing 3D Consistent Human Avatars with Sparse inputs via Gaussian Splatting and contrastivE learning, as shown in Fig. 1. We first integrate a skeleton-driven rigid deformation and a non-rigid cloth dynamics deformation to create a human avatar. To enhance 3D consistency under sparse inputs, we utilize the intrinsic 3D consistency of images across different poses within the same person. Specifically, for each training pose/image, we select a similar pose/image from the dataset and then adjust the deformed Gaussians using the proposed Dynamic Avatar Adjustment (DAA), an explicit point-based control graph adjustment strategy, to the selected similar pose. Then, we minimize differences between the rendered image of the adjusted Gaussians and the image corresponding to the selected similar pose, which serves as an additional form of supervision for human avatars. Additionally, we employ 3D geometry contrastive learning, utilizing features from a 3D feature extractor, to further enhance the global 3D consistency of generated human avatars. We conduct extensive experiments on the ZJU-MoCap data [26], H36M [10] and surprisingly find that CHASE outperforms other SOTAs in both full and sparse inputs setting. Our work makes the following contributions:
@@
  • We propose a 3D geometry contrastive learning to enforce consistency across different representations of the same pose and enhance global 3D understanding.
@@
2 Related Work
2.1 Contrastive Representation Learning
@@
Contrastive Representation Learning is one of the mainstream self-supervised learning paradigms, which learns potential semantics from constructed invariance or equivariance. In 3D, PointContrast [39] proposes geometric augmentation to generate positive and negative pairs. CrossPoint [1] uses both inter- and intra-modal contrastive learning. PointCLIP [45] achieves image-point alignment by projecting point clouds onto 2D depth images. RECON [28] focuses on single- and cross-modal contrastive learning through discriminative contrast [13] or global feature alignment [30]. Our CHASE introduces a 3D geometry contrastive learning method to enforce consistency across different representations of the same pose.
2.2 3D Editing and Deformation
@@
Traditional deformation methods in computer graphics are typically based on Laplacian coordinates [4], Poisson equations [43], and cage-based methods [41]. However, these methods often rely on implicit and computationally expensive NeRF-based approaches.
@@
Numerous works [3, 47] have proposed techniques for editing 3D Gaussian Splatting (3DGS)[12]. SuGaR[5] introduces a mesh extraction method that produces meshes from 3DGS, which can then be edited. SC-GS [9] proposes deforming Gaussians by transferring the movement of control points. Our CHASE employs a novel explicit point-based control graph deformation strategy, which is more efficient than previous methods.
2.3 3D Human Modeling
@@
Since the high-quality rendering achieved by the seminal work Neural Radiance Fields (NeRF)[21], there has been a surge of research on neural rendering for human avatars[18, 17, 26, 31]. Although NeRF is designed for static objects, HumanNeRF [38] extends NeRF to enable capturing dynamic human motion using just a single monocular video. Neural Body [26] associates a latent code with each SMPL [19] vertex to encode appearance, which is then transformed into observation space based on the human pose. Furthermore, Neural Actor [18] learns a deformable radiance field with SMPL [19] as guidance and utilizes a texture map to improve the final rendering quality. Posevocab [17] designs joint-structured pose embeddings to encode dynamic appearances under different key poses, allowing for more effective learning of joint-related appearances. However, a major limitation of NeRF-based methods is that NeRFs are slow to train and render.
@@
Point-based rendering [48] has proven to be an efficient alternative to NeRFs for fast inference and training. Extending point clouds to 3D Gaussians, 3D Gaussian Splatting (3DGS)[12] models the rendering process by splatting a set of 3D Gaussians onto the image plane via alpha blending. Given the impressive performance of 3DGS in both quality and speed, numerous works have further explored the 3D Gaussian representation for dynamic 3D human avatar reconstruction[16, 29, 8, 14]. Human Gaussian Splatting [22] showcases 3DGS as an efficient alternative to NeRF. SplattingAvatar [32] and GomAvatar [37] extend lifted optimization to simultaneously optimize the parameters of the Gaussians while walking on the triangle mesh. However, these methods struggle to maintain 3D consistency and produce low-quality reconstructions when applied to human avatar creation with only sparse inputs. Our CHASE introduces a novel 2D image supervision to 3D human body modeling and 3D geometry contrastive learning, enhancing the 3D consistency of human avatars.
@@
3 Preliminaries
SMPL [19]. The SMPL model is a pre-trained parametric human model representing body shape and pose. In SMPL, body shape and pose are controlled by pose and shape. In this work, we apply the Linear Blend Skinning (LBS) algorithm used in SMPL to transform points from a canonical space to a posed space.
LBS [33]. Linear Blend Skinning (LBS) is a weight-based technique that associates each vertex with one or more joints and uses weight values to describe the influence of each joint on the vertex. Vertex deformation is calculated by linearly interpolating transformations on the associated joints: $\mathcal{X}_{v}^{{}^{\prime}}=\sum_{j=1}^{J}w_{j}(\mathcal{X}_{v})B_{j}\mathcal{X}_{v}$, where $J$ represents the number of joints, $N$ represents the number of vertices, $\mathcal{X}_{v}^{{}^{\prime}}\in\mathbb{R}^{N\times 3}$ is the new position of the skinned vertex, $w\in\mathbb{R}^{N\times J}$ is the skinning weight matrix, $B\in\mathbb{R}^{J\times 4\times 4}$ is the affine transformation matrix of each joint representing rotation and translation, (i.e. bone transforms) and $\mathcal{X}_{v}\in\mathbb{R}^{N\times 3}$ is the original mesh vertex position.
3D Gaussian Splatting (3DGS) [12]. 3DGS explicitly represents scenes using point clouds, where each point is modeled as a 3D Gaussian defined by a covariance matrix $\Sigma$ and a center point $\mathcal{X}$, the latter referred to as the mean. The value at point $\mathcal{X}$ is:
@@ 3 Preliminaries
In rendering novel views, differential splatting as introduced by [42], involves using a viewing transform $W$ and the Jacobian matrix $J$ of the affine approximation of the projective transformation to compute the transformed covariance matrix: $\Sigma^{\prime}=JW\Sigma W^{T}J^{T}$. The color and opacity at each pixel are computed from the Gaussian’s representations: $G(\mathcal{X})=e^{-\frac{1}{2}\mathcal{X}^{T}\Sigma^{-1}\mathcal{X}}.$ The blending of $N$ ordered points overlapping a pixel is given by the formula: $\mathcal{C}=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{i}),$ where $c_{i}$, $\alpha_{i}$ represent the density and color of this point computed by a 3D Gaussian $G$ with covariance $\Sigma$ multiplied by an optimizable per-point opacity and SH color coefficients.
4 Method
@@
4.1 Non-rigid and Rigid Deformation
Inspired by [38, 29], we deform 3D Gaussians from canonical space ${\mathcal{G}_{c}}$ to observation space ${\mathcal{G}_{o}}$ by integrating a rigid articulation with a non-rigid transformation. We employ a non-rigid deformation network that takes the canonical positions $\mathcal{X}_{c}$ of the 3D Gaussians ${\mathcal{G}c}$ and a pose latent code which encodes SMPL pose $p_{i}$ using a lightweight hierarchical pose encoder [20]. The network then outputs the offsets for various parameters of the 3D Gaussians $\mathcal{G}_{c}$: $\Delta{(\mathcal{X},\mathcal{C},\alpha,s,r)}$. The canonical Gaussians are deformed by:
@@ 4.1 Non-rigid and Rigid Deformation
Figure 3: For each training pose/image, we select similar poses/images from the dataset and then adjust the deformed Gaussians using DAA. By minimizing the difference between the rendered image of the adjusted avatar and the selected similar pose image, we introduce additional supervision, thereby refining the creation of photo-realistic and animatable avatars.
4.2 Dynamic Avatar Adjustment
To address extremely sparse inputs, we leverage the intrinsic 3D consistency of human avatars across different poses/images, as shown in Fig. 3. Though the same pose may exhibit slight variations in non-rigid deformation, these differences occupy only a small number of pixels from a whole-body perspective and thus have minimal impact on the overall body reconstruction, which we further demonstrate in the experimental Section. 5.3.
Specifically, for each training pose/image, we select a similar pose $p^{a}_{i}$ with its paired image $x_{i}^{a}$ by computing the orientation and limb angle difference provided by SMPL [19] model from the dataset. Then we use a dense motion field $F_{adj}$ as an additional adjustment to transform deformed Gaussians ${\mathcal{G}_{o}}$ into adjusted Gaussians ${\mathcal{G}_{a}}$, aligning them with the selected pose/image ($p_{i}^{a}$/$x_{i}^{a}$). In this way, we successfully introduce an additional 2D image supervision, improving the 3D consistency of human avatars.
To achieve precise control of the 3D Gaussians, we sample 6,890 points from the SMPL model [19] as our sparse control points in canonical space. Then, we obtain the dense motion field using LBS by locally inheriting the LBS weights from neighboring control points. Specifically, for each 3D Gaussian, we use the k-nearest neighbor (KNN) search to find its nearest neighboring control points in canonical space. The entire adjustment process is as: